# Transformation of orientation and rotation angles of synchronous satellites: Application to the Galilean moons

### v1

@@ 2.3 Projection of the spin axis in the reference plane
Following the habits of the rotation modeling community (e.g. Bills [2005], Baland et al. [2012]), we introduce the components ($n_{x},n_{y},n_{z}$) of the unit vector along the orbit normal expressed in the coordinates of the Laplace reference frame:
@@ 3.2 Approximate transformations
With the exception of $s_{z}$, each angle and unit vector component is expressed as the sum of the Laplace pole value, a slow linear term with a rate noted $tr$, and a series noted $\Delta$ of periodic terms with small amplitudes:
@@
where $t$ is the time. The Laplace pole values of $(s_{x},s_{y},s_{z})$ are by definition $(0,0,1)$. The values of $\alpha_{LP}$ and $\delta_{LP}$ depend on the chosen orbital model, see Section 2. The periodic terms $\Delta s_{x}$ and $\Delta s_{y}$ are the nutations in space of the spin axis (the dominant part of the nutational motion can be called precession). The periodic parts $\Delta\alpha_{S}(t)$ and $\Delta\delta_{S}(t)$ can also be called nutations. In the present context of approximate transformations, all these periodic variations are considered as first order terms. By construction, the spin models for $s_{x}$ and $s_{y}$ also contain secular terms. Since for the Galilean satellites they are much smaller than the largest periodic terms, these secular terms are considered to be of second order. Therefore, since the spin vector is unitary, $\Delta s_{z}\simeq-\left(\Delta s_{x}^{2}+\Delta s_{y}^{2}\right)/2$ and is formally considered here as a second-order quantity. By construction, $s_{z}$ should also include quadratic and Poisson terms. Very small quadratic and Poisson terms (product between a linear term and a periodic series) are neglected here in $s_{x/y}$ and $\alpha_{S}/\delta_{S}$, because they are formally considered of order 3 or above and do not significantly improve accuracy in numerical applications.
To obtain approximate expressions for $\Delta\alpha_{S}$ and $\Delta\delta_{S}$, assuming that we have expressions for $\Delta s_{x}$ and $\Delta s_{y}$, we express the exact trigonometric expressions of Eqs. (5) correct-up to the second-order in $\Delta$ quantities. Extracting $\Delta\alpha_{S}$ and $\Delta\delta_{S}$, we find:
@@
The opposite transformation is given by:
@@
The transformation coefficients in Eqs. (8-9) depend on the Laplace plane orientation and slightly differ for each Galilean satellite. Numerical values of the first order coefficients for the four Galilean satellites are listed in Table 1. The second order transformation coefficients are usually between 1-1 and $5$. Similar expressions, but restricted to the first-order part of the transformation, apply to the small trends ($tr$ quantities):
@@
The first order series in $\Delta\alpha_{s}$ and $\Delta\delta_{S}$ (Eqs. 8 without the second order terms) keep an analytical/trigonometrical form similar as that of $s_{x},s_{y}$ (that is to say, series with the same number of terms, at the same frequencies). The second-order terms result in additional periodic terms to the series, due to the recombination of sine and cosine factors. This recombination may introduce a small constant term in addition to the periodic series. Note that the distinction between first and second order is purely formal, and not based on the actual values of the amplitudes. Some of these second-order terms may be of larger amplitude than some of the first-order terms. For this reason, the second-order terms allow to improve the accuracy of the transformation (from $0.01^{\circ}$ to less than $0.0001^{\circ}$ in $\alpha$, and from $0.006^{\circ}$ to $0.0001^{\circ}$ in $\delta$), see Table 10 and Section 6.3.
@@ 3.3 Back to the obliquity and node longitude
We use the notation $\arctan{(x,y)}$ in the definitions for $\psi$ and $\Omega$ to take into account the quadrant in which the point $(x,y)$ lies. If we take into account the quadrant in which the projections of $\boldsymbol{\hat{\textbf{s}}}$ and $\boldsymbol{\hat{\textbf{n}}}$ onto the Laplace plane lie when calculating $\psi$ and $\Omega$, then $\theta$ and $i$ will be positive by default. This is convenient for satellites whose spin and orbit poles are on the same side with respect to the Laplace pole. For a satellite like the Moon, whose spin axis and orbit normal are located on opposite sides with respect to the Laplace pole, and thus their projections in opposite quadrants, it can be interesting from a modeling point of view (where $\psi$ is generally assumed to be equal to or close to $\Omega$) to change the sign of $\theta$ and add $\pi$ to $\psi$. This is equivalent to redefining the descending node as the ascending node, but with a negative obliquity.
@@
$\psi$ $\approx$ $\arctan{(-\Delta s_{y},\Delta s_{x})}+\frac{(tr_{y}\,\Delta s_{x}-tr_{x}\,\Delta s_{y})}{\Delta s_{x}^{2}+\Delta s_{y}^{2}}\,t$
$\approx$ $\arctan{(-\Delta\delta_{S},\cos\delta_{LP}\,\Delta\alpha_{S})}+\frac{1}{2}\Delta\alpha_{S}\sin\delta_{LP}\left(\frac{\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}+2\Delta\delta_{S}^{2}}{\Delta\delta_{S}^{2}+\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}}\right)+\frac{(tr_{\delta}\,\Delta\alpha_{S}-tr_{\alpha}\,\Delta\delta_{S})\cos\delta_{LP}}{\Delta\delta_{S}^{2}+\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}}\,t.$
Due to the linear terms in $s_{x}$ and $s_{y}$ (or in $\alpha_{S}$ and $\delta_{S}$), these relations contain some terms proportional to the linear rates $tr_{x/y/\alpha/\delta}$ that we here keep since they ensure a higher degree of precision (e.g. error $\leq 0.0001^{\circ}$ for Europa’s inertial obliquity compared to the solution obtained with Eq. (11a), instead of $0.002^{\circ}$ without Poisson terms). For Io, those terms contribute up to $6^{\circ}$ in $\psi$ after $100$ years, see section 6.2.
As mentioned in the introduction to this section, the transformations of Section 3.2 also apply to $\boldsymbol{\hat{\textbf{n}}}=(n_{x},n_{y},n_{z})$ and $(\alpha_{O},\delta_{O})$. Therefore similar relationships as Eqs. (14) exist for the inclination $i$ and node longitude $\Omega$, with $\Delta n_{x}$ and $\Delta n_{y}$ instead of $\Delta s_{x}$ and $\Delta s_{y}$ and $\Delta\alpha_{O}$ and $\Delta\delta_{O}$ instead of $\Delta\alpha_{S}$ and $\Delta\delta_{S}$. For the same reason, this also allows us to find a workable expression for the orbital obliquity $\varepsilon$ :
$\varepsilon$ $\approx$ $\sqrt{(\Delta s_{x}-\Delta n_{x})^{2}+(\Delta s_{y}-\Delta n_{x})^{2}},$ (15a)
$\approx$ $\sqrt{\cos\delta_{LP}^{2}(\Delta\alpha_{S}-\Delta\alpha_{O})^{2}+(\Delta\delta_{S}-\Delta\delta_{O})^{2}}-\frac{(\Delta\alpha_{O}-\Delta\alpha_{S})^{2}(\Delta\delta_{O}+\Delta\delta_{S})\sin 2\delta_{LP}}{4\sqrt{\cos\delta_{LP}^{2}(\Delta\alpha_{S}-\Delta\alpha_{O})^{2}+(\Delta\delta_{S}-\Delta\delta_{O})^{2}}}.$ (15b)
@@ 3.4 Offset with respect to the Cassini plane
If the satellite’s orbit and spin precess regularly at a constant rate and if the satellite is unaffected by tidal dissipation, then the Laplace pole $\boldsymbol{\hat{\textbf{l}}}$, the orbit normal $\boldsymbol{\hat{\textbf{n}}}$ and the spin axis $\boldsymbol{\hat{\textbf{s}}}$ belong to the same plane, known as the Cassini plane. The Cassini plane is defined as the plane containing both the Laplace pole and the orbit normal. Due to irregular orbital precession and/or coupling with polar motion and/or tidal dissipation, the spin axis is in fact almost never in the Cassini plane. The offset $\zeta$ of the spin axis with respect to the Cassini plane can be written as, see Yseboodt and Margot [2006], Baland et al. [2017]:
@@
With this sign convention, the offset is defined as a prograde angle. In the case of a dissipative body with a regular retrograde orbital precession, a positive offset indicates a spin axis lagging ”behind” the Cassini plane.
@@
$\zeta$ $\simeq$ $\frac{\Delta n_{x}\,\Delta s_{y}-\Delta n_{y}\,\Delta s_{x}}{\sqrt{\Delta n_{x}^{2}+\Delta n_{y}^{2}}}+\frac{\left(\Delta n_{x}\,(\Delta n_{x}-\Delta s_{x})+\Delta n_{y}\,(\Delta n_{y}-\Delta s_{y})\right)}{\left(\Delta n_{x}^{2}+\Delta n_{y}^{2}\right)^{3/2}}(tr_{y}\,\Delta n_{x}-tr_{x}\,\Delta n_{y})\,t,$ (17b)
@@ 4 Transformation between the rotation angles $\phi$ and $W$
The difference between the rotation angles $W$ and $\phi_{Euler}$ is the angle $\beta$ corresponding to the angle $\Omega_{ISLS}$ in Fig. 1:
@@ 4.2 Approximate transformation
Using the approximate expressions for $s_{x}$ and $s_{y}$ of Eqs. (7a-7b) and the transformation relationship between the different sets of orientation angles, we now find approximations for $\beta$, either as a function of $tr_{x}$, $\Delta s_{x}$ and $\Delta s_{y}$, or as a function of $tr_{\alpha}$, $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ (we neglect quadratic and Poisson terms):
@@
From these expressions, we can see that, although the two angles do not belong to the same plane, $\beta$ and $\psi$ have fairly close values at all times (maximum difference of around $0.7^{\circ}$ for Ganymede). They both exhibit a precession (linear trend) which causes them to deviate relatively rapidly from their epoch values. In the following, we define the difference between $\beta$ and $\psi$ as the angle $\mu$, that includes the difference between the trends and periodic variations of the two angles:
@@
4.3 The different definitions of $\phi$, in relation to the different definitions of the libration angle $\gamma$
Librations in longitude are defined as the periodic variations of the satellite’s rotation around its spin axis. We have just defined two different rotation angles ($\phi_{Euler}$ and $W$), whose periodic variations $\Delta\phi_{Euler}$ and $\Delta W$ are different from each other, since $\beta$ and varies periodically with time. Clearly, if the definition of the rotation angle is not unique, neither is that of libration. The rotation modeling community still uses other definitions, based on dynamic considerations.
To resolve for the rotation variations with respect to synchronicity with the orbital motion, the gravitational torque in libration models depends on the difference between a rotation angle and the orbital true longitude (see Section 5.2), with both angles measured from the same origin. The choice of a fixed/inertial origin is necessary to correctly describe the torque and find a realistic libration motion. The most elementary libration model considers that the satellite evolves on a non-precessing eccentric Keplerian orbit (which can therefore be assimilated, for modeling purposes only, to both a reference plane instead of the Laplace plane and to the satellite’s equator) and that it undergoes diurnal libration such that the satellite’s long axis points towards the central body at pericenter (e.g. Van Hoolst et al. 2008). In this context, it makes sense to measure the angle of rotation on the plane of the orbit, starting from the pericenter. We note that rotation angle $\phi_{pericenter}$. It is also customary to refer the satellite’s orbit not to its Laplace plane, but to Jupiter’s equatorial plane, since in this modeling context, the Laplace plane is not of paramount importance. The geometric relationship between $\phi_{Euler}$ and $\phi_{pericenter}$, at first order in small angles $i$, $\theta$ and $\varepsilon$ is
@@
where both the left and right hand sides are the broken angle starting at the point $IQ$ in Fig. 1 and ending at the meridian plane, whereas the $\phi$ angles start at different points. $\Omega_{JupEq}$ is the node longitude of the orbit starting from the intersection between the ICRF equator and Jupiter Equator at J2000 and $\omega_{JupEq}$ is the argument of the satellite pericenter with respect to Jupiter Equator. $\omega_{JupEq}+\pi$ is the argument of the pericenter of the planet around the satellite. The $\Omega_{IQ\,IL}$ angle is constant, since it is defined by three inertial planes. For a Keplerian non-precessing orbit, $\psi$, $\Omega_{JupEq}$ and $\omega_{JupEq}$ can be considered constant as well, so that $\Delta\phi_{Euler}=\Delta\phi_{pericenter}$ and is essentially diurnal.
@@
The points IL and QL could also be suitable, since the triangle defined by the points IQ, QL, and IL is fixed. Depending on the actual starting point (IQ, IL or QL), $\phi_{Inertial}$ is defined to within one constant, just as $\Omega_{JupEq}$.
@@
where $M(t)$ only varies linearly at the rate $n_{M}$ and $\gamma$ is proportional to the eccentricity $e$ and $\sin M$, so that the satellite’s major axis always points approximately towards the orbit’s empty focus, and the angle of rotation is the sum of an epoch value, of a linear trend representative of synchronous rotation, and of a single periodic term. For more advanced modeling, a similar formulation should be used, where $M(t)$ is replaced by the mean mean longitude $L$ or by the mean longitude $\mathcal{L}$ (following here the notations in Lainey et al. [2006]), both measured along Jupiter’s equator:
@@
If the orbit is close to singular (this is the case for the Galilean satellites), these trends may be difficult to fit separately.
@@
With the latter expressions, the meaning of $\mu$ becomes clearer. If $\beta$ represents the transformation between $W$ and $\phi_{Euler}$, then $\mu$ represents, to within one constant, the transformation between $W$ and $L+\gamma_{u}$ (or equivalently between $W$ and $\mathcal{L}+\gamma_{f}$), or more simply, between $W$ and $\phi_{inertial}$:

### v2

@@ 2.3 Projection of the spin axis in the reference plane
Following previous rotation studies (e.g. Bills [2005], Baland et al. [2012]), we introduce the components ($n_{x},n_{y},n_{z}$) of the unit vector along the orbit normal expressed in the coordinates of the Laplace reference frame:
@@ 3.2 Approximate transformations
With the exception of $s_{z}$, each angle and unit vector component is expressed as the sum of the values of the Laplace pole coordinates, a slow linear term with a rate noted $tr$, and a series noted $\Delta$ of periodic terms with small amplitudes:
@@
where $t$ is the time. The Laplace pole location in the coordinates of the frame attached to the Laplace frame is by definition given by $(0,0,1)$. The values of $\alpha_{LP}$ and $\delta_{LP}$ depend on the chosen orbital model, see Section 2. The periodic terms $\Delta s_{x}$ and $\Delta s_{y}$ are the nutations in space of the spin axis (the dominant part of the nutational motion, if it exists, can be called precession by analogy with Earth and Mars). The periodic parts $\Delta\alpha_{S}(t)$ and $\Delta\delta_{S}(t)$ can also be called nutations. In the present context of approximate transformations, all these periodic variations are considered as first order terms. By construction, the spin models for $s_{x}$ and $s_{y}$ also contain secular terms. Since for the Galilean satellites, over a 100-year interval around J2000 (the time interval we target for the spin model), they are much smaller than the largest periodic terms, these secular terms are considered to be of second order. Therefore, since the spin vector is a unit vector, $\Delta s_{z}\simeq-\left(\Delta s_{x}^{2}+\Delta s_{y}^{2}\right)/2$ and is formally considered here as a second-order quantity. By construction, $s_{z}$ should also include quadratic and Poisson terms. Very small quadratic and Poisson terms (product between a linear term and a periodic series) are also neglected here in $\alpha_{S}/\delta_{S}$, because they are formally considered of order 3 or above and do not significantly improve accuracy in numerical applications.
To obtain approximate expressions for $\Delta\alpha_{S}$ and $\Delta\delta_{S}$, assuming that we have expressions for $\Delta s_{x}$ and $\Delta s_{y}$, we express the exact trigonometric expressions of Eqs. (6) correct up to the second-order in $\Delta$ quantities. Extracting $\Delta\alpha_{S}$ and $\Delta\delta_{S}$, we find, see also Lieske [1993a, b]:
@@
The inverse transformation is given by:
@@
The transformation coefficients in Eqs. (9-10) depend on the Laplace plane orientation and slightly differ for each Galilean satellite. Numerical values of the first and second-order coefficients for the four Galilean satellites are listed in Table 1. Similar expressions, but restricted to the first-order part of the transformation, apply to the small trends ($tr$ quantities):
@@
The first-order series in $\Delta\alpha_{S}$ and $\Delta\delta_{S}$, obtained from Eqs. (9) by neglecting second-order terms, have an analytical trigonometric form similar to that of $s_{x},s_{y}$. In particular, they involve the same number of terms and the same set of frequencies. The second-order terms result in additional periodic terms to the series, due to the recombination of sine and cosine factors. This recombination may introduce a small constant term in addition to the periodic series. Note that the distinction between first and second order is purely formal, and not based on the actual values of the amplitudes. Some of these second-order terms may be of larger amplitude than some of the first-order terms. For this reason, the second-order terms allow to improve the accuracy of the transformation (from $0.01^{\circ}$ to less than $0.0001^{\circ}$ in $\alpha$, and from $0.006^{\circ}$ to $0.0001^{\circ}$ in $\delta$), see Table 10 and Section 6.3.
@@ 3.3 Back to the obliquity and node longitude
Following the notation of the built-in Mathematica ArcTan function, we use the notation $\arctan{(x,y)}$ in the definitions for $\psi$ and $\Omega$ to take into account the quadrant in which the point $(x,y)$ lies. If we take into account the quadrant in which the projections of $\boldsymbol{\hat{\textbf{s}}}$ and $\boldsymbol{\hat{\textbf{n}}}$ onto the Laplace plane lie when calculating $\psi$ and $\Omega$, then $\theta$ and $i$ will be positive by default. This is convenient for satellites whose spin and orbit poles are on the same side with respect to the Laplace pole. For a satellite like the Moon, whose spin axis and orbit normal are located on opposite sides with respect to the Laplace pole, and thus their projections in opposite quadrants, it can be useful from a modeling point of view (where $\psi$ is generally assumed to be equal to or close to $\Omega$) to change the sign of $\theta$ and add $\pi$ to $\psi$. This is equivalent to redefining the descending node as the ascending node, but with a negative obliquity.
@@
$\psi$ $\approx$ $\arctan{(-\Delta s_{y},\Delta s_{x})}+\frac{(tr_{y}\,\Delta s_{x}-tr_{x}\,\Delta s_{y})}{\Delta s_{x}^{2}+\Delta s_{y}^{2}}\,t,$
$\approx$ $\arctan{(-\Delta\delta_{S},\cos\delta_{LP}\,\Delta\alpha_{S})}+\frac{1}{2}\Delta\alpha_{S}\sin\delta_{LP}\left(\frac{\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}+2\Delta\delta_{S}^{2}}{\Delta\delta_{S}^{2}+\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}}\right)$
$+\frac{(tr_{\delta}\,\Delta\alpha_{S}-tr_{\alpha}\,\Delta\delta_{S})\cos\delta_{LP}}{\Delta\delta_{S}^{2}+\Delta\alpha_{S}^{2}\cos^{2}\delta_{LP}}\,t.$
Due to the linear terms in $s_{x}$ and $s_{y}$ (or in $\alpha_{S}$ and $\delta_{S}$), these relations contain some terms proportional to the linear rates $tr_{x/y/\alpha/\delta}$, or Poisson terms, which we retain since they ensure a higher degree of precision. For example, the resulting error on Europa’s inertial obliquity with respect to Eq. (12a) is $\leq 0.0001^{\circ}$ when these terms are included, compared to $0.002^{\circ}$ when the Poisson terms are omitted. For Io, those terms contribute up to $6^{\circ}$ in $\psi$ after $100$ years, see section 6.2.
As mentioned in the introduction to this section, the transformations of Section 3.2 also apply to $\boldsymbol{\hat{\textbf{n}}}=(n_{x},n_{y},n_{z})$ and $(\alpha_{O},\delta_{O})$. Therefore similar relationships to Eqs. (15) exist for the inclination $i$ and node longitude $\Omega$, with $\Delta n_{x}$ and $\Delta n_{y}$ instead of $\Delta s_{x}$ and $\Delta s_{y}$ and $\Delta\alpha_{O}$ and $\Delta\delta_{O}$ instead of $\Delta\alpha_{S}$ and $\Delta\delta_{S}$. For the same reason, this also allows us to find a workable expression for the orbital obliquity $\varepsilon$:
$\varepsilon$ $\approx$ $\sqrt{(\Delta s_{x}-\Delta n_{x})^{2}+(\Delta s_{y}-\Delta n_{x})^{2}},$ (16b)
$\approx$ $\sqrt{\cos\delta_{LP}^{2}(\Delta\alpha_{S}-\Delta\alpha_{O})^{2}+(\Delta\delta_{S}-\Delta\delta_{O})^{2}}$
$-\frac{(\Delta\alpha_{O}-\Delta\alpha_{S})^{2}(\Delta\delta_{O}+\Delta\delta_{S})\sin 2\delta_{LP}}{4\sqrt{\cos\delta_{LP}^{2}(\Delta\alpha_{S}-\Delta\alpha_{O})^{2}+(\Delta\delta_{S}-\Delta\delta_{O})^{2}}}.$
@@ 3.4 Offset with respect to the Cassini plane
If the satellite’s orbit and spin precess regularly at a constant rate and if the satellite is unaffected by tidal dissipation, then the Laplace pole $\boldsymbol{\hat{\textbf{l}}}$ (the normal to the Laplace plane), the orbit normal $\boldsymbol{\hat{\textbf{n}}}$ and the spin axis $\boldsymbol{\hat{\textbf{s}}}$ belong to the same plane, known as the Cassini plane. The Cassini plane is defined as the plane containing both the Laplace pole and the orbit normal. Due to quasi-periodic orbital precession and/or coupling with polar motion and/or tidal dissipation, the spin axis is in fact almost never in the Cassini plane. The offset $\zeta$ of the spin axis with respect to the Cassini plane can be written as, see Yseboodt and Margot [2006], Baland et al. [2017]:
@@
With this sign convention, the offset is defined as a prograde angle. In the case of a dissipative body with a regular retrograde orbital precession, a positive offset indicates a spin axis lagging “behind” the Cassini plane.
@@
$\zeta$ $\simeq$ $\frac{\Delta n_{x}\,\Delta s_{y}-\Delta n_{y}\,\Delta s_{x}}{\sqrt{\Delta n_{x}^{2}+\Delta n_{y}^{2}}}$ (18b)
$+\frac{\left(\Delta n_{x}\,(\Delta n_{x}-\Delta s_{x})+\Delta n_{y}\,(\Delta n_{y}-\Delta s_{y})\right)}{\left(\Delta n_{x}^{2}+\Delta n_{y}^{2}\right)^{3/2}}(tr_{y}\,\Delta n_{x}-tr_{x}\,\Delta n_{y})\,t,$
@@ 4 Transformation between the rotation angles $\phi$ and $W$
In this manuscript, $\dot{\xi}$ denotes the linear rate of the angle $\xi$, while $\xi^{\prime}$ is used to denote time derivatives.
The difference between the rotation angles $W$ and $\phi_{Euler}$ is the angle $\beta$ corresponding to the angle $\Omega_{IS\,LS}$ (ISIS is the intersection of the equator on the ICRF plane) in Fig. 1:
@@ 4.2 Approximate transformation
Using the approximate expressions for $s_{x}$ and $s_{y}$ of Eqs. (8a-8b) and the transformation relationship between the different sets of orientation angles, we now find approximations for $\beta$, either as a function of $tr_{x}$, $\Delta s_{x}$ and $\Delta s_{y}$, or as a function of $tr_{\alpha}$, $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ (we neglect quadratic and Poisson terms):
@@
From these expressions, we can see that, although the two angles do not belong to the same plane, $\beta$ and $\psi$ have fairly close values at all times (maximum difference of around $0.7^{\circ}$ for Ganymede). They both exhibit a precession (linear trend) which causes them to deviate relatively rapidly from their epoch values. In the following, we define the difference between $\beta$ and $\psi$ as the angle $\mu$, that includes the difference between the trends and periodic variations of the two angles (see also Lieske [1993a, b] for $\Delta\mu$):
@@
4.3 The periodic variations in $W$, including the physical librations
Librations in longitude are defined as the periodic variations of the satellite’s rotation around its spin axis. We have just defined two different rotation angles ($\phi_{Euler}$ and $W$), whose periodic variations $\Delta\phi_{Euler}$ and $\Delta W$ are different from each other, since $\beta$ varies periodically with time. Clearly, if the definition of the rotation angle is not unique, neither is that of libration.
To resolve for the rotation variations with respect to synchronicity with the orbital motion, the gravitational torque in libration models depends on the difference between a rotation angle and the orbital true longitude (see Section 5.2), with both angles measured from the same origin. The choice of a fixed/inertial origin is necessary to correctly describe the torque and find a realistic libration motion. The most elementary libration model considers that the satellite evolves on a non-precessing eccentric Keplerian orbit and that it undergoes diurnal libration such that the satellite’s long axis points towards the central body at pericenter (e.g. Van Hoolst et al. 2008). For modeling purposes only, this Keplerian orbit can therefore be used both as the reference plane (instead of the Laplace plane) and as a proxy for the satellite’s equatorial plane. In this context, it makes sense to measure the angle of rotation on the plane of the orbit, starting from the pericenter. We note that rotation angle $\phi_{pericenter}$. It is also customary to refer the satellite’s orbit not to its Laplace plane, but to Jupiter’s equatorial plane, since in this modeling context, the Laplace plane is not of paramount importance. The geometric relationship between $\phi_{Euler}$ and $\phi_{pericenter}$, at first order in small angles $i$, $\theta$ and $\varepsilon$ is
@@ 4.3 The periodic variations in $W$, including the physical librations
where both the left and right hand sides are the broken angle starting at the point $IQ$ (the intersection of Jupiter equator on the ICRF plane) in Fig. 1 and ending at the meridian plane, whereas the $\phi$ angles start at different points. $\Omega_{JupEq}$ is the node longitude of the orbit starting from the intersection between the ICRF equator and Jupiter Equator at J2000 and $\omega_{JupEq}$ is the argument of the satellite pericenter with respect to Jupiter Equator. $\omega_{JupEq}+\pi$ is the argument of the pericenter of the planet around the satellite. The $\Omega_{IQ\,IL}$ angle is constant, since it is defined by three inertial planes. For a Keplerian non-precessing orbit, $\psi$, $\Omega_{JupEq}$ and $\omega_{JupEq}$ can be considered constant as well, so that $\Delta\phi_{Euler}=\Delta\phi_{pericenter}$ and is essentially diurnal.
@@
The points $IL$ and $QL$ could also be suitable, since the triangle defined by the points $IQ$, QLQL, and $IL$ is fixed. Depending on the actual starting point ($IQ$, ILIL or $QL$), $\phi_{Inertial}$ is defined to within one constant, just as $\Omega_{JupEq}$.
@@
where $M(t)$ only varies linearly at the rate $n_{M}$ and $\gamma$ is proportional to the eccentricity $e$ and $\sin M$, so that the satellite’s major axis always points approximately towards the orbit’s empty focus, and the angle of rotation is the sum of an epoch value, of a linear trend representative of synchronous rotation, and of a single periodic term. For more advanced modeling, a similar formulation should be used, where $M(t)$ is replaced by the mean mean longitude $L$ or by the mean longitude $\mathcal{L}$ (following here the notations in Lainey et al. 2006), both measured along Jupiter’s equator:
@@
If the orbit is nearly circular and has a low inclination (this is the case for the Galilean satellites), these trends may be difficult to fit separately.
@@
With the latter expressions, the meaning of $\mu$ becomes clearer. If $\beta$ represents the transformation between $W$ and $\phi_{Euler}$, then $\mu$ represents, to within one additive constant, the transformation between $W$ and $L+\gamma_{u}$ (or equivalently between $W$ and $\mathcal{L}+\gamma_{f}$), or more simply, between $W$ and $\phi_{inertial}$: